A pitcher throws a 0.145 kg baseball at 40.19 m/s at the catcher (who is standing still). It takes the catcher 0.07 seconds to stop the baseball after catching it. What was average force required to catch the ball?
Answer #1

F =dp/dt = 0.145*40.19/0.07 = 83.25 N (ans)